### v2

@@ Abstract
Particle detectors based on scintillators are widely used in high-energy physics and astroparticle physics experiments, nuclear medicine imaging, industrial and environmental detection, etc. Precisely extracting scintillation signal characteristics at the event level is important for these applications, not only in respect of understanding the scintillator itself, but also kinds and physical property of incident particles. Recent researches demonstrate data-driven neural networks surpass traditional statistical methods, especially when the analytical form of signals is hard to obtain, or noise is significant. However, most densely connected or convolution-based networks fail to fully exploit the spectral and temporal structure of scintillation signals, leaving large space for performance improvement. In this paper, we propose a network architecture specially tailored for scintillation pulse characterization based on previous works on time series analysis. The core insight is that, by directly applying Fast Fourier Transform on original signals and utilizing different frequency components, the proposed network architecture can serve as a lightweight and enhanced representation learning backbone. We prove our idea in two case studies: (a) simulation data generated with the setting of the LUX dark matter detector, and (b) experimental electrical signals with fast electronics to emulate scintillation variations for the NICA/MPD calorimeter. The proposed model achieves significantly better results than the reference model in literature and densely connected models, and demonstrates higher cost-efficiency than conventional machine learning methods.
@@ 1 Introduction
Particle detectors based on scintillators are widely used in high-energy physics and astroparticle physics experiments [1], nuclear medicine imaging [2, 3], industrial and environmental detection [4, 5], etc. The scintillation light is converted to electrical signals by transducers like photomultipliers (PMT) or silicon photomultipliers, and then processed by subsequent circuits. With the rapid development of electronic devices and integrated circuits, nowadays high bandwidth front-end electronics and high sampling rate analog-to-digital converters become affordable even for medium or large detector systems. This generates a considerable amount of raw data coming out of digitizers. The data explosion in nuclear instrumentation offers opportunities as well as challenges for the community, which gives rise to researches on software algorithms [6, 7] and related online hardware implementation [8, 9, 10] for signal processing.
@@
In this work, we demonstrate that neural network architectures inspired by spectral analysis of scintillation signals can be a justified choice for multiple regression and classification tasks. This network architecture originates from TimesNet [32], a recent advance in time series analysis, but some important adjustments are incorporated to make the model lightweight and enhanced in terms of performance. We call the new model TimesNet-LE. Rather than making a horizontal comparison between different network architectures, this work aims at demonstrating how the performance of simple networks can be boosted with representation learning implemented by the TimesNet-LE backbone. The major innovations and contributions of the paper include:
@@ 3.1.1 Simulation setup
In should be noted that, under the above conditions, using traditional analyzing methods like curve fitting is inefficient because of the complicated effects of scintillation emission, optical process, electronics and noise. Neural networks become a natural choice if sufficient data are generated.
@@ 3.1.2 Results
In simulation, the 10-GSPS raw waveform is downsampled with a ratio 20:1, and a section of 256 samples is randomly selected on the downsampled waveform. This gives an input time series with 2-ns period (or equivalent 500-MSPS rate) and 256 elements. We have generated 10000 ER events and 10000 NR events, forming a dataset with 20000 examples in total. Then the dataset is divided into a training set with 14000 examples, a validation set with 2000 examples and a test set with 4000 examples.
@@
Figure 5: Performance measures (standard deviation of residuals between the ground-truth and the predicted) of parameter estimation on simulation data with correlated or uncorrelated noise at different signal-to-noise ratios. Network architectures with TimesNet-LE are marked with stars.
@@ 3.2.1 Experiment setup
Figure 8: Diagram and photograph of the experimental system.
@@
In the experiment, we use the fast electronics to emulate variations of scintillation signals. The fast electronics is taken from a shashlik electromagnetic calorimeter [36] (with alternate plastic scintillators and lead absorbers) in the Nuclotron-based Ion Collider fAcility (NICA) experiment [37]. The detector system used in the experiment is shown in Fig. 8. Square pulses with preset amplitudes of voltage are generated by the arbitrary function signal generator (Tektronix AFG3102C). These pulses are injected into the fast electronics through an adapter board for power & test injection. The electronics is surrounded by a shielding box to avoid electromagnetic interference. Test pulses reach the single-channel front ends with preset capacitance to produce response signals. Then the response signals are wired to another adapter board for readout and connected to a custom application-specific integrated circuit (WRX2F04A1 with 1-GSPS rate and 14-bit precision) for analog-to-digital conversion. Digitized signals are finally sent to computers for storage and analysis.
@@ 3.2.2 Results
In the experiment, the 1-GSPS raw waveform is downsampled with a ratio 5:1, and a section of 384 samples is randomly selected on the downsampled waveform. This gives an input time series with 5-ns period (or equivalent 200-MSPS rate) and 384 elements. We have gathered approximately 4143 events for each condition with specific injection voltage and front-end capacitance, forming a dataset with 99438 examples in total. Then the dataset is divided into a training set with 69600 examples, a validation set with 9942 examples and a test set with 19896 examples.
@@
Finally, we select random forest and two best-performing network architectures and plot the confusion matrices for classifying the cross conditions of injection voltage and front-end capacitance. These are shown in Fig. 12, for applying random forest, Fig. 13, for applying the linear head on the TimesNet backbone, and Fig. 14, for applying the nonlinear head on the TimesNet-LE backbone. For random forest, there are light-colored squares scattering in a wide area in Fig. 12, which shows the errors are random and not restricted to a few conditions. For the linear head on the TimesNet backbone, there are several light-colored lines parallel to the main diagonal line in Fig. 13, which corresponds to the misclassified examples into the wrong classes of the front-end capacitance, and this is the main source of decreased overall accuracy for this network architecture. In contrast, the confusion matrix in Fig. 14 is the cleanest with few misclassified examples in the upper left region, and almost all examples fall into the main diagonal elements. The detailed visualization demonstrate the fitness of TimesNet-LE as a representation learning backbone for multi-class classification.
@@ 4 Discussion
It should be noted that features from the same network architecture have the potential to serve some more sophisticated tasks, like uncertainty estimation [39] or covariance matrix estimation [40], if possible with the signal quality and sampling rate.
@@ Appendix A Details of simulation
where $\mathscr{F}[\cdot]$ and $\mathscr{F}^{-1}[\cdot]$ denote Fourier and inverse Fourier transform. Then the covariance matrix of the noise is generated by sampling the auto-correlation function. Finally, the correlated noise is sampled from a multivariate Gaussian distribution with zero mean and this covariance matrix.

### v3

@@ Abstract
Particle detectors based on scintillators are widely used in high-energy physics and astroparticle physics experiments, nuclear medicine imaging, industrial and environmental detection, etc. Precisely extracting scintillation signal characteristics at the event level is important for these applications, not only in respect of understanding the scintillator itself, but also kinds and physical property of incident particles. Recent researches demonstrate data-driven neural networks surpass traditional statistical methods, especially when the analytical form of signals is hard to obtain, or noise is significant. However, most densely connected or convolution-based networks fail to fully exploit the spectral and temporal structure of scintillation signals, leaving large space for performance improvement. In this paper, we propose a network architecture specially tailored for scintillation pulse characterization based on previous works on time series analysis. The core insight is that, by directly applying Fast Fourier Transform on original signals and utilizing different frequency components, the proposed network architecture can serve as a lightweight and enhanced representation learning backbone. We prove our idea in two case studies: (a) simulation data generated with the setting of the LUX dark matter detector, and (b) experimental electrical signals with fast electronics to emulate scintillation variations for the NICA/MPD calorimeter. The proposed model achieves significantly better results than the reference model in literature and densely connected models and demonstrates higher cost-efficiency than conventional machine learning methods.
@@ 1 Introduction
Particle detectors based on scintillators are widely used in high-energy physics and astroparticle physics experiments [1], nuclear medicine imaging [2, 3], industrial and environmental detection [4, 5], etc. The scintillation light is converted to electrical signals by transducers like photomultipliers (PMT) or silicon photomultipliers and then processed by subsequent circuits. With the rapid development of electronic devices and integrated circuits, nowadays high bandwidth front-end electronics and high sampling rate analog-to-digital converters become affordable even for medium or large detector systems. This generates a considerable amount of raw data coming out of digitizers. The data explosion in nuclear instrumentation offers opportunities as well as challenges for the community, which gives rise to researches on software algorithms [6, 7] and related online hardware implementation [8, 9, 10] for signal processing.
@@
In this work, we demonstrate that the neural network architecture inspired by spectral analysis of scintillation signals can be a justified choice for multiple regression and classification tasks. This network architecture originates from TimesNet [32], a recent advance in time series analysis, but some important adjustments are incorporated to make the model lightweight and enhanced in terms of performance. We call the new model TimesNet-LE. Rather than making a horizontal comparison between different network architectures, this work aims at demonstrating how the performance of simple networks can be boosted with representation learning implemented by the TimesNet-LE backbone. The major innovations and contributions of the paper include:
@@ 3.1.1 Simulation setup
In should be noted that, under the above conditions, using traditional analyzing methods like curve fitting is inefficient because of the complicated effects of scintillation emission (equation 1), optical process (equation 2), electronics (equation 3) and noise (equation 5). Directly combining these factors will result in a frequency-domain system function with 6 poles and several zeros. The locations of poles and zeros will change with input parameters, which can be quite ill-conditioned when solving the reverse optimization problem. Besides, considering noise correlation will incur a large covariance matrix in curve fitting, which is very compute-intensive. On the other hand, neural networks become a natural choice if sufficient data are generated.
@@ 3.1.2 Results
In simulation, the 10-giga samples per second (GSPS) raw waveform is downsampled with a ratio 20:1, and a section of 256 samples is randomly selected on the downsampled waveform. This gives an input time series with the 2-ns period, or equivalently the 500-mega samples per second (MSPS) rate, and 256 elements. We have generated 10,000 ER events and 10,000 NR events, forming a dataset with 20,000 examples in total. Then the dataset is divided into a training set with 14,000 examples, a validation set with 2000 examples and a test set with 4000 examples.
@@
Figure 5: Performance measures (standard deviation of residuals between the ground-truth and the predicted) of parameter estimation on simulation data with correlated or uncorrelated noise at different signal-to-noise ratios. Network architectures with TimesNet-LE are marked with stars. Sub-figures in the bottom row share the same legend with those in the top row.
@@ 3.2.1 Experiment setup
Figure 8: Diagram and photograph of the experimental system. (HV: High Voltage, ADC: Analog-to-Digital Converter, GSPS: Giga Samples Per Second)
@@
In the experiment, we use the fast electronics to emulate variations of scintillation signals. The fast electronics is taken from a shashlik electromagnetic calorimeter [36] (with alternate plastic scintillators and lead absorbers) for the Multi-Purpose Detector (MPD) in the Nuclotron-based Ion Collider fAcility (NICA) [37]. The detector system used in the experiment is shown in Fig. 8. Square pulses with preset amplitudes of voltage are generated by the arbitrary function signal generator (Tektronix AFG3102C). These pulses are injected into the fast electronics through an adapter board for power & test injection. The electronics is surrounded by a shielding box to avoid electromagnetic interference. Test pulses reach the single-channel front ends with preset capacitance to produce response signals. Then the response signals are wired to another adapter board for readout and connected to a custom application-specific integrated circuit (WRX2F04A1 with 1-GSPS rate and 14-bit precision) for analog-to-digital conversion. Digitized signals are finally sent to computers for storage and analysis.
@@ 3.2.2 Results
In the experiment, the 1-GSPS raw waveform is downsampled with a ratio 5:1, and a section of 384 samples is randomly selected on the downsampled waveform. This gives an input time series with 5-ns period (or equivalent 200-MSPS rate) and 384 elements. We have gathered approximately 4143 events for each condition with specific injection voltage and front-end capacitance, forming a dataset with 99,438 examples in total. Then the dataset is divided into a training set with 69,600 examples, a validation set with 9942 examples and a test set with 19,896 examples.
@@
Finally, we select random forest and two best-performing network architectures and plot the confusion matrices for classifying the cross conditions of injection voltage and front-end capacitance. These are shown in Fig. 12 (for applying random forest), Fig. 13 (for applying the linear head on the TimesNet backbone), and Fig. 14 (for applying the nonlinear head on the TimesNet-LE backbone). For random forest, there are light-colored squares scattering in a wide area in Fig. 12, which shows the errors are random and not restricted to a few conditions. For the linear head on the TimesNet backbone, there are several light-colored lines parallel to the main diagonal line in Fig. 13, which corresponds to the misclassified examples into the wrong classes of the front-end capacitance, and this is the main source of decreased overall accuracy for this network architecture. In contrast, the confusion matrix in Fig. 14 is the cleanest with few misclassified examples in the upper left region, and almost all examples fall into the main diagonal elements. The detailed visualization demonstrate the fitness of TimesNet-LE as a representation learning backbone for multi-class classification.
@@ 4 Discussion
It should be noted that the proposed method is generally applicable to nuclear radiation detectors, not restricted to scintillation detectors. For example, in pulse sampling-based fluorescence lifetime imaging [39], TimesNet-LE can be used to estimate fluorescence lifetime values from waveforms coming out of avalanche photodiodes or microchannel plate photomultiplier tubes. In three-dimensional position-sensitive detection [40], TimesNet-LE can be used for multi-channel feature extraction with semiconductor-coupled silicon detectors. In data acquisition systems of high-energy physics experiments [41], TimesNet-LE can be used for on-detector data pre-processing with double-sided silicon strip detectors.
Finally, features from the same network architecture have the potential to serve some more sophisticated tasks, like uncertainty estimation [42] or covariance matrix estimation [43], if possible with the signal quality and sampling rate.
@@ Acknowledgements
Declarations
[…]
Competing Interests
The authors declare that they have no competing interests.
@@ Appendix A Details of simulation
where $\mathscr{F}[\cdot]$ and $\mathscr{F}^{-1}[\cdot]$ denote Fourier and inverse Fourier transforms. Then the covariance matrix of the noise is generated by sampling the auto-correlation function. Finally, the correlated noise is sampled from a multivariate Gaussian distribution with zero mean and this covariance matrix.